If 6.00mL of 0.1000 M NaOH solution is needed to just neutralize excess acid after 20.00 mL of 0.1000 M HCl was added to 1.00 g of an antacid, how many moles of acid can the antacid counteract per gram?
Answer #1

NaOH will neutralise = 6* 0.1=0.6 *10^-3 moles


so antacid will neutralise=(2-0.6) *10^-3 =1.4 *10^-3 moles